Name: Prefix-1

Description:

Prefix-1 is a command to access an auxiliary keymap in the Emacs tradition. There are 3 such keymaps: the default keymap, the Prefix-1 keymap and the Prefix-2 keymap. Prefix-1 is always bound to <ESC>. Normally, the user need not be aware of this command; it simply exists more commands can be run from the keyboard. You can think of it as a different kind of shift key.

See Also: Prefix-2

Name: Prefix-2

Description:

Prefix-2 is the command to access the second auxiliary keymap. Prefix-2 is always bound to <C-X>.

See Also: Prefix-1

Name: ansi-keys

Description:

This function interprets the codes sent by the Sun function keys and attempts to do something intelligent with them. The best solution would be to be able to map each of them to a function just like any other key. But since the function keys generate strings of characters, it isn’t easy to incorporate them into dvitool’s key binding scheme. So instead dvitool interprets each sequence and runs a command from a private key map, much like the key map the menu uses. The person who installed dvitool can change the default bindings for the function keys; the normal defaults are “unbound” for every key except the arrow keys on the right keypad which scroll, L5 which runs toggle-tool, and L7 which runs close-window. These binding are site-wide; once compiled in by the installer, they cannot be changed by individual users. This command is something of a hack; it was never intended to be run intentionally by the user. It’s ability to interpret the ANSI sequences depends on it being bound to \e[ and you will after you type some additional input, you will get innocuous error messages if you run ansi-keys manually. For all its short comings though, it does suppress the unwanted characters that result when your fingers slip onto one of the function keys.

See Also: enable-ansi-keys Name: ascii-of-selection

Description:

This function displays a previously-made selection in ASCII to make it easier to determine what the correct search string for the selection would be.

See Also: erase-selection, extend-selection, select-char

Name: beginning-of-file

Description:

This command positions dvitool on the first page of the DVI file.

See Also: end-of-file

Name: bind-to-key

Usage: function-name key-strokes

Description:

This command “binds” a key (or keys) to a function that will be executed when that key is pressed. The binding’s life is the life of the dvitool that executed it; to get bindings which take effect on every instance of dvitool, put the bindto-key command in your startup file. Control and mouse characters have ASCII representations which must be used in the startup file. Describe-key displays what command a particular key sequence will invoke; as a side effect, it also displays the ASCII representation of the sequence, so it can be used to easily determine the correct ASCII representation for any valid input.

See Also: describe-key, dump-bindings, Startup-File

Name: bottom-edge

Description:

This command positions the page image so the bottom edge of the current DVI page is visible. If a numeric argument is given, the bottom edge of the visible page (without the bottom margin) will be shown.

See Also: bottom-left, bottom-right, top-edge Name: bottom-left Description:

This command positions the DVI page so the bottom left hand corner is visible. If a numeric argument is given, the bottom left corner of the visible page (without the margins) will be shown.

See Also: bottom-edge, bottom-right, top-left

Name: bottom-right Description:

This command positions the DVI page so the bottom right hand corner is visible.

See Also: bottom-edge, bottom-left, top-right

Name: bound-to Usage: command name Description:

This command describes all of the key strokes that will invoke command name separated by commas.

See Also: dump-bindings

Name: cd Usage: filename Description:

This command is similiar to the sh(1) or csh(1) commands of the same name; it changes the directory. When it is given a null argument, it changes to the user’s home directory. The working directory can be viewed with the print command; it can also be altered with the set command, but since cd takes a file name which can be completed on, and cwd is considered a string when changed with set, it is generally easier to use cd.

See Also: cwd, print, set

Name: close-window Description:

This command “closes” dvitool to a small iconic shape. The image that is painted and the position at which the image is painted are controlled by user-definable variables.

See Also: init-icon-file, init-icon-x, init-icon-y Name: describe-key

Usage: key-strokes

Description:

This command identifies what command a sequence of keystrokes will invoke. In the process it also echos the ASCII character sequences dvitool uses to represent all of the different input combinations. This means that you don’t have to remember that control–shift–middle–mouse button is represented as \m\^M; describe-key will tell you that.

See Also: bind-to-key, dump-bindings

Name: dump-bindings

Description:

This command creates a file in the current directory called dvitool.bindings that describes which command each of the possible keyboard and mouse combinations will invoke.

See Also: bind-to-key, dump-commands

Name: dump-commands

Description:

This command creates a file in the current directory called dvitool.commands that describes each of the dvitool commands, the type of arguments that it takes, and the key strokes it is bound to separated by commas. In addition, all of the dvitool variables and the their types are described.

See Also: dump-bindings

Name: end-of-file

Description:

This command positions dvitool on the last page of the DVI file.

See Also: beginning-of-file Name: erase-selection Description:

This command erases the current selection, if any. It is primarily used to ensure that a search operation begins at the top of the page. A search begins at the end of the current selection, if any, or the top of the page, so erasing the current selection guarantees that the search will begin at the top of the current page.

See Also: select-char

Name: exec Usage: command name Description:

Exec is the most general way to run a dvitool command. Exec takes a (possibly completed) string and looks in it’s table of command names. If some command name exactly matches the string, then that command is executed. Since the command name argument to exec is subject to completion, a quick way to see all of dvitool’s commands is to type <ESC>x?.

Name: exit Description:

This is the command to use when you wish to end your session with dvitool. It is usually bound to e, <C-C>, and <C-X><C-C>.

Name: exit-help Description:

This command exits from the last instance of help to the document you were viewing before you invoked help. If the stack of help entries is empty, an error message is printed.

See Also: help-commands, help-overview, help-variables

Name: expose-tool Description:

This command behaves exactly the same as the “expose” entry on a the standard tool menu; i.e. it causes a partially covered dvitool to become the topmost tool, and thus unobscured. If dvitool is not covered by some other tool, this command is a no-op.

See Also: hide-tool, toggle-tool Name: extend-selection Description:

This command extends a previously made selection to include new characters. The command select-char only selects 1 character; thus, the usual method to selection a region of characters is to first select one boundary character with select-char and then to select the region by selecting the other boundary character with extendselection. It is an error to run this command with no current selection.

See Also: ascii-of-selection, erase-selection, select-char

Name: find-file Usage: filename Description:

This command is the main entry point for viewing a DVI file. Find-file takes a filename, expands ~ and $ characters into home directories and environment variables respectively, appends .dvi if necessary and attempts to read the file and display it. Various error messages are issued if the named file is not a DVI file or if it malformed etc. The argument filename is subject to completion, so at any time the user may enter a space character to attempt completion or a question mark to see the list of choices dvitool will select from. If a question mark or space character is needed in the filename itself, precede it with a backslash.

In addition to completion, there is “next file” and “previous file” selection, used in conjuction with the DVI files named on the command line. Control-N gets the next file name off of the list and control-P gets the previous file name. So to preview all of the chapters of your upcoming book, invoke dvitool like this: %␣dvitool␣ch*.dvi and then use control-N and control-P to select which file you’d like to view. Just as with space and question mark, if you really want a control-N or a control-P in your filename, precede it with a backslash.

Name: full-screen Description:

The full-screen command toggles the size of dvitool between as big as the physical screen will allow and whatever size it was before full-screen was invoked.

See Also: zoom-horizontal, zoom-vertical

Name: goto-manuscript-page Usage: string Description:

This command searches for a specific page by examining between 1 and 10 of TEX’s \count variables. Each page of a DVI file contains the value of these 10 variables when it was shipped out; with the proper macro definitions, any number of section, chapter, or heading configurations can be described with them. The string that goto-manuscript-page takes is parsed into up to 10 fields; then each of the \count fields of each page is compared with the parsed string until either a match is found or until all the pages have been examined. The format for the search string is n[.n] … where n is either a decimal number or the asterisk character *. The dot character is a field separator. Thus, assuming the the appropriate macros have put the chapter number into \count1, the string *.4 will find the first page of chapter 4 beyond the current page.

Goto-manuscript-page also accepts strings of roman numerals as input.

See Also: goto-physical-page

Name: goto-physical-page Usage: integer Description:

This command provides an alternative way to goto-manuscript-page to seek to a DVI page. Instead of the complexity of \count variables, goto-physical-page simply displays the nth page of the file. Thus, goto-physical-page with an argument of 1 is semantically equivalent to beginning-of-file.

See Also: goto-manuscript-page

Name: help-commands Usage: help command name Description:

This command displays the dvitool command help file and then searchs for the optional argument in that file. The argument is subject to completion. If completion is used to generate the search string, a match is guaranteed. The user may type in an arbitrary string as an argument.

See Also: help-overview, help-variables

Name: help-overview Usage: help overview string Description:

This command displays the overview file which describes the general mechanisms of dvitool. The optional string argument describes a search string to search for in the overview file. If completion is used to generate the string, a match is guaranteed; however, the user may search for arbitrary strings in the overview file.

See Also: help-commands, help-variables

Name: help-variables Usage: help variable name Description:

Like help-commands, this command displays the dvitool variable help file and then searchs for an optional string in that file. If completion is used to generate the search string, then a match is guaranteed; the user may supply arbitrary strings though.

See Also: help-commands, help-overview

Name: hide-tool Description:

This command behaves exactly the same as the “Hide” option of the standard tool menu; i.e. it causes dvitool to be placed at the bottom of the stack of visible tools.

See Also: expose-tool, toggle-tool

Name: left-edge Description:

This command positions the page image so the left edge of the page is visible. If a numeric argument is given, the left edge of the visible page (without the margin) will be shown. This command is commonly used when one wants to see the maximum amount of visible page.

See Also: bottom-left, right-edge, top-left

Name: list-all-commands Description:

This command simply displays the names of all of the commands.

See Also: dump-commands, list-all-variables Name: list-all-variables Description:

This command simply displays the names of all of the variables.

See Also: dump-commands, list-all-commands

Name: magstep-0 Description:

This function displays the DVI file at its normal size. It is a no-op unless the file has previous been displayed at some other magstep. Note that magnification in dvitool is global, that is, everything including the width of the page and the margins will be affected by the magnification routines. There is no analogy to TEX’s true points

      in dvitool.

See Also: magstep-minus-5…magstep-5

Name: magstep-1 Description:

This command magnifies the document to 120 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5

Name: magstep-2 Description:

This command magnifies the document to 144 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5

Name: magstep-3 Description:

This command magnifies the document to 172.8 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5

Name: magstep-4 Description:

This command magnifies the document to 207.4 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5 Name: magstep-5 Description:

This command magnifies the document to 248.8 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5

Name: magstep-minus-1 Description:

This command shrinks the document to 83.3 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5

Name: magstep-minus-2 Description:

This command shrinks the document to 69.4 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5

Name: magstep-minus-3 Description:

This command shrinks the document to 57.9 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5

Name: magstep-minus-4 Description:

This command shrinks the document to 48.2 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5

Name: magstep-minus-5 Description:

This command shrinks the document to 40.2 percent of its magstep-0 size.

See Also: magstep-minus-5…magstep-5 Name: mouse-menus Description:

This command invokes the menus. It may only be executed in response to a mouse button, though bind-to-key will bind it anywhere.

Name: next-page Description:

This is the normal command to page through the DVI file. It displays the next page of the document with the upper left hand corner of the page visible, easing the normal reading flow from the bottom of one page to the top of the next.

See Also: goto-manuscript-page, next-page-positioned, previous-page

Name: next-page-positioned Description:

This command displays the next page of the DVI file at the same position on the page as the current page. Thus, if you are viewing the bottom of one page and want to see the bottom of the next page, this command suffices.

See Also: next-page, previous-page-positioned

Name: numeric-argument Description:

This command is used to enter a “count” which can be interpreted by functions that take an integer as their first argument. It is a prefix argument that is zeroed after each command. The command interpreter checks each time a command is run for a non-zero value of the numeric argument. If it is non-zero and the first argument of the command being executed is of type integer, the command interpreter substitutes the numeric argument for the command argument. For example, goto-physical-page takes a single integer argument and is usually bound to G. So the character sequence 1G will position dvitool on the first page of the DVI file. Other commands use the presence or absence of a numeric argument as a switch; for example, the edge commands act differently when a numeric argument is present.

Name: previous-page Description:

This is the previous analog to next-page.

See Also: next-page, previous-page-positioned Name: previous-page-positioned Description:

This command is the previous analog to next-page-positioned.

See Also: next-page-positioned, previous-page

Name: print Usage: variable name Description:

This function shows the value of dvitool’s internal variables.

See Also: set

Name: redraw-tool Description:

This command behaves exactly the same as the “Redisplay” option of the standard tool menu; i.e. it causes dvitool to repaint all of its windows.

Name: reload-fonts Description:

This function is identical to reread-file except that before the file is reread, all of the characters in the font cache are flushed. This command is most useful when some of the pixel images of characters in a DVI file have changed.

See Also: reread-file

Name: reread-file Description:

This command reloads and redisplays the current DVI file. It is most useful when in the edit–TEX–preview cycle.

See Also: find-file

Name: right-edge Description:

This command positions the page image so the right edge of the current DVI page is visible. If a numeric argument is given, the right edge of the visible page (without the margin) will be shown. This command is commonly used when one wants to see the maximum amount of visible page.

See Also: bottom-right, left-edge, top-right Name: scroll-absolute Usage: integer Description:

This command scrolls down vertically down the page by the amount of the argument. The argument is a percentage between 0 and 100 of the page to scroll down.

See Also: scroll-down

Name: scroll-down Description:

This command scrolls the page image down in the window if there is any more of the page to display vertically. The amount scrolled by default is 1/3 of the vertical size of the window; the default can be changed by supplying a numeric argument. That argument is taken as the denominator of the fraction of the window to scroll down. For example, the keystrokes 10d will scroll down by 1/10 the vertical size of the window.

See Also: numeric-arg, scroll-left, scroll-right, scroll-up

Name: scroll-left Description:

This command is the leftward analog of scroll-down.

See Also: scroll-down

Name: scroll-right Description:

This command is the rightward analog of scroll-down.

See Also: scroll-down

Name: scroll-up Description:

This command is the upward analog of scroll-down.

See Also: scroll-down Name: search-backward Usage: literal-string Description:

This command is exactly like search-forward except that it searches backwards through the document instead of forwards, and the search is begun either from the last character of the page, or the character before the first character of the selection.

See Also: search-backward-by-font, search-forward

Name: search-backward-by-font Usage: font-name literal-string Description:

This command is exactly like search-backward except that a SET CHAR must be in the named font to be a match.

See Also: search-backward, search-forward

Name: search-forward Usage: literal-string Description:

This command searches forward sequentially from a point on the current page for a string of characters in the DVI file. The string is first parsed for any control sequences (any sequence of characters beginning with the backslash (`\') character.) Then the search begins, either at the first character on the page if there is no selection on the current page or at the first character after the selection. Each character in the search string is compared with a SET CHAR operand in the DVI file or with a dvitool-created logical space character that matches an ASCII space character. The character matches if the integer parameter to the SET CHAR matches the integer value of the ASCII character. If the end of the search string is reached, a match is reported. Note that a match which spans a page boundary will not be found. The variables kern-threshold and line-break-threshold control how dvitool decides whether movements are interword spaces or kerns.

Since a match is based on the position of the character in the font, ligatures, in particular require some special handling. The “fi” ligature for example is at position 12 decimal in the roman font family. This corresponds to the ASCII character control-L. So to search for the word “file”, you must use the string \^Lle. The command ascii-of-selection greatly eases the task of computing the proper search string for a selection.

Searches can can be aborted prematurely by hitting the L1 key (usually the key at the upper-left most position on the keyboard). This is the same key that is used to reboot the machine, but you needn’t hold it down like a control key to make it work; just depress it and release it. Every so often dvitool will check to see if the key has been depressed and will terminate the search if it has. This key is effective only in the search functions.

See Also: ascii-of-selection, kern-threshold, line-break-threshold, search-backward, searchforward-by-font

Name: search-forward-by-font Usage: font-name literal-string Description:

This command is exactly like search-forward except that a SET CHAR must be in the named font to be a match.

Dvitool keeps a cache of all the fonts you’ve used, so the fonts presented as choices for the font-name argument may or may not actually be in your DVI file. This isn’t a major concern, however, since dvitool will just issue an innocuous error message if you ask to restrict searching to a font that doesn’t exist in your current DVI file.

See Also: search-forward

Name: select-char Description:

This command makes a single character of the DVI file the current selection. The character chosen is highlighted by displaying it in inverse video. It may only be executed in response to an input from the mouse. To extend the selection to more than one character, use extend-selection. A “double click” will select a word (a sequence of contiguous non-space characters).

See Also: ascii-of-selection, extend-selection, which-char, which-font

Name: set Usage: variable name value string Description:

This command is used to change the value of dvitool’s variables. It can be run interactively, but is used most often in the user’s .dvitoolrc file to customize dvitool’s behavior. This first argument is the name of the variable to change and the second argument is the value to change it to. The second argument will of course vary depending on the type of the variable to be changed; to get the previous contents of the variable, type your rprnt character (usually control-R, see stty(1)).

See Also: print

Name: toggle-tool Description:

This command puts dvitool at the top of the windows visible on the screen unless it already is the top window; in that case it puts it on the bottom of the stack.

See Also: expose-tool, hide-tool

Name: top-edge Description:

This command positions the page image so the top edge of the current DVI page is visible. If a numeric argument is given, the top edge of the visible page (without the top margin) will be shown.

See Also: top-left, top-right, bottom-edge

Name: top-left Description:

This command positions the DVI page so the top left hand corner is visible. If a numeric argument is given, the top left corner of the visible page (without the margins) will be shown.

See Also: top-edge, top-right, bottom-left

Name: top-right Description:

This command positions the DVI page so the top right hand corner is visible. If a numeric argument is given, the top right corner of the visible page (without the margins) will be shown.

See Also: top-edge, top-left, bottom-right

Name: version Description:

Print a string which contains the version number of dvitool.

Name: which-char Description:

This command describes some of the properties of the first character of the selection. It reports the font name and the position of the character in the font in decimal, octal, and hexadecimal. If the selection contains more than one character, it is trimmed down to the first charcter in the selection and the information is reported on that character.

See Also: ascii-of-selection, which-font

Name: which-font Description:

This command reports the font name of the characters in the selection. If there are characters from more that one font in the selection, the selection is first trimmed to include only characters from the same font as the first character in the selection.

See Also: ascii-of-selection, which-char

Name: zoom-horizontal Description:

Make dvitool as wide as the physical screen. Like full-screen and zoom-vertical, this command is a toggle; running it again will return dvitool to its previous dimensions.

See Also: full-screen, zoom-vertical

Name: zoom-vertical Description:

This command offers the same functionality as clicking control middle button on the name stripe, e.g. the tool becomes a tall as the physical screen while retaining the current width. This command is a toggle; running it again will return dvitool to its previous dimensions.

See Also: full-screen, zoom-horizontal